A board 170 feet long is cut into two pieces such that the length of one piece is two-thirds of the length of the other piece. Find the length of the shorter piece of board
Answer by rfer(16322)

one piece is two thirds the other so we have five intervals
170/5=34
3*34=102ft
2*34=68ft